OpenStudy (anonymous):

What is a possible first step to eliminate the variable x in the following system? 2x – 7y = 3 4x + 5y = –2 Multiply the first equation by 7. Multiply the first equation by –2. Multiply the second equation by 2. Multiply the second equation by –3.

OpenStudy (anonymous):

c

OpenStudy (anonymous):

thank you !

OpenStudy (radar):

Careful there Nana3456, look at the second choice also.

OpenStudy (radar):

Take the first equation and multiply it by 2 -2 (2x-7y = 3) -4x +14y = -6 now add to the 2nd equation 4x + 5y = -2 Getting: ---------------- 19y = -8 or y= -8/19 Now you can substitute that value for y and solve for x. The point for you to note, is x could be eliminated by multiplying the first equation by -2.

OpenStudy (anonymous):

Thank you !